A committee of 4 has to be selected from amongst 5 men and 4 women. In how many ways can this be done so as to include at least 1 man and 1 woman.
Found 2 solutions by Jk22, robertb:
Answer by Jk22(389)

Choosing 1 man and 1 woman gives 5*4 possibilities,

remain to choose 2 people out of 4+3=7 : 7*6

Since the order does not influence, this makes :

5*4*7*6/4! = 5*4*7*6/(4*3*2) = 35 possibilities

Answer by robertb(5830)

# ways of choosing 1 man from 5 = 5C1 = 5.
# ways of choosing 1 woman from 4 = 4C1 = 4.
# ways of choosing 2 people from 7 remaining people = 7C2 = 21.
By the fundamental principle of counting, the #ways of choosing 4 people so as to include at least 1 man and 1 woman is 5*4*21 = 420.